k y quest6h BH yöu you may may have. Please be sure to answer all questions the best you can. have. Please be sure to answer all questions the best you can Par L (d Polntsy Thlnking Critically. Q1. (3 points) Consider each of the scenarios: a. Can -0.41 be the probability of some event? Why or why not? b. Can 1.29 be the probability of some event? Why or why not? c. Can 0.86 be the probability of some event? Why or why not? Q2. (1 points) Suppose todays weather forecast on weather.com says there is a 40% probability of rain today. What is the compliment of the event rain today? What is the probability of that compliment? Part II.(5 points) Basic Probability Use the following for Q1 and Q2. Suppose in a class of 140 students 35 of them are Biology Majors. Q1. (I point) If a student is chosen at random what is the likelihood they will be a Biology Major? P(Biology Major)-
Answer #1

1)

(a) No, there can not be a negative probability of any event as the probability of any event lies in the range of [0,1]

(b) No, the probability of any event can not be greater than 1.

(c) Yes, the probability of 0.86 is possible as it lies in the range of [0,1]

2)

The compliment of the event "rain today" is "no rain today"

We know that the sum of the probabilities of an event and it's compliment is equal to 1.

Thus, probability of the compliment, no rain today = 1 - 0.4 = 0.6

3)

P(Biology major) = No. of Biology majors/Total no. of students

= 35/140 = 0.25